Shawniqua is 8 years younger than HiKiem. Their combined ages are 58.
How old is Shawniqua?
How old is HiKiem?

Answer by Cromlix(4381)

Hi there,
Make HiKiem's age = x
so, Shawniqua = x - 8
Their combined ages are 58.
x + x - 8 = 58
Collect like terms
2x = 58 + 8
2x = 66
x = 33.
HiKiem is 33 years old
Shawniqua is 25 years old.
Hope this helps :-)